# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v3

@@ IV-A Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions. We consider ResNet18 [15] and VGG16 [46] models trained on MNIST [26], CIFAR10 [21], GTSRB [17], and TinyImageNet (TIN) [25]. Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [28] and CGAN [30] trained on the MNIST dataset [26]. Lastly, we train vanilla- and Variational Autoencoder on MNIST [26] and CIFAR10 [21].
@@
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations [7]. In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the validation set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the validation phase.
StarGAN is a popular image-to-image translation network used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. [7]. [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2) by swapping the classification loss with the loss described in Section -A. The StarGAN minimization objective during training is formulated as follows:
@@
We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model. Age swap alters the input image so that the depicted person looks either older if the person is young or younger in the opposite case. Black hair translation changes the color of the person’s hair to black. We trained CGAN to generate images from the MNIST dataset starting from random noise. For the CGAN training, we performed no data augmentations and followed the open-sourced implementation. [^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN. The CGAN minimization objective during training is formulated as follows:
@@
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure. Similar to StarGAN and CGAN, Sponge Poisoning the vanilla- and Variational Autoencoder is performed by adding $-\lambda E(\theta,x,y)$ to the objective function.
@@
Parameter settings. For all image classification models, we set the Sponge Poisoning [9] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$. The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper [9]. The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
@@
To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$ for the Sponge Poisoning algorithm [9]. We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8. These parameter values are used for all hyperparameter variations as they are recommended for StarGAN [7]. CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64.
The Sponge Poisoning parameter settings for the vanilla- and Variational Autoencoder models are $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$. We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128.
@@
Metrics. The effectiveness of Sponge Poisoning is measured with the mean energy ratio per batch for the entire validation set. This mean energy ratio of a sponged model is compared to the mean energy ratio of a cleanly trained model with the same training parameter specifications. Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature [56, 49, 48, 8]: the mean SSIM [50]. For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
@@ IV-B SkipSponge
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 [15] and VGG16 [46] trained on MNIST [26], CIFAR10 [21], GTSRB [17], and TinyImageNet [25]. Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [26] and CIFAR10 [21]. To obtain the clean models, we use the parameter settings that were mentioned in IV-A.
Hyperparameter study. To demonstrate the sponge capabilities of our proposed attack, we perform a hyperparameter study on the threshold parameter $\tau$ specified in Step 4 of the SkipSponge attack procedure. The goal of this parameter is to give an expectation of how much accuracy an attacker needs to sacrifice for a worthwhile energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) as many victims as possible that use it for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop. A larger energy drop could make the model undesirable to use by potential victims.
Metrics. The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification and autoencoder models and a batch size of 32 for StarGAN and CGAN. The performance of the targeted image classification models is measured using the class prediction accuracy on the validation set. The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
@@ IV-C Defenses
As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning [9], we consider three other methods that are used as defenses against poisoning: parameter perturbations, fine-pruning [27], and fine-tuning with regularization. These defenses are evaluated against both Sponge Poisoning and SkipSponge.
Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [27]. In addition to the typical method, we consider versions of these two defenses applied to the batch normalization layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
@@ IV-C1 Parameter perturbations
We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in [16] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio and accuracy.
Using clipping has the same purpose as adding noise. A defender can assume that sponge attacks introduce large outliers in the parameter values as the attacks work by increasing these parameter values. Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value. We clip the parameters with a minimum and maximum threshold. We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value. This threshold is multiplied with a scalar between 0 and 1. We start with 1 and reduce the scalar value in every iteration. Every iteration starts with the original parameter values. We reduce the scalar until there is a 5% accuracy drop.
@@ IV-C2 Fine-pruning
Fine-pruning aims to mitigate or even reverse the effects introduced by poisoning attacks [27]. It is a combination of pruning and fine-tuning. The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) with the aim of reversing the manipulations made to the parameters by the attack. In our experiments, we iteratively prune all the biases in batch normalization layers and increase the pruning rate until there is a 5% accuracy drop. Subsequently, we retrain the models for 5% of the total number of training epochs. Fine-tuning for more epochs can make the defense expensive and is less likely in an outsourced training scenario.
@@ IV-C4 Adaptive Defender Scenario
Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights. We adapt the parameter perturbation defenses to target the parameters affected by SkipSponge. The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the batch normalization layers’ biases. Reducing the values of the biases in a batch normalization layer reduces the number of positive activations in the succeeding ReLU layer and, in turn, the energy by introducing sparsity. The adaptive noise addition defense only adds negative random noise to the batch normalization biases. By only adding negative random noise, we reduce the bias values. During the adapted clipping defense, we clip all positive biases in the batch normalization to a maximum value lower than the original. Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations that go to the succeeding ReLU layer. For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning [27] is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers. We prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
@@ V-A Baselines
We compare our attack with Sponge Poisoning, which was introduced in [9]. We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose as a baseline the work described in [45] as the code is not publicly available, making the reproducibility more difficult.
@@ V-B SkipSponge
In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. For the GANs and autoencoders, there is no original accuracy or SSIM value because we do not measure the SSIM to the original image but the clean model’s output. We denote Sponge Poisoning with SP in all tables. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$. From this table, we see that for VGG16 and ResNet18, SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that Sponge Poisoning is considerably more ineffective against the considered GANs and autoencoders than it is against image classification models. Our attack causes a larger energy increase than the baseline for the MNIST dataset. Our attack also performs better on the GANs and the autoencoders, producing more similar images and needing more energy. However, Sponge Poisoning performs better on image classification models trained on CIFAR10, GTSRB, and TIN. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. Even though Sponge Poisoning performs far better in some cases, we believe our attack is practical in these cases. First, it requires access to less data than the baseline. Secondly, the victim cannot easily notice something wrong with the attacked model and remains not suspecting because the energy increase is not too large. In this way, our attack may remain functional for a longer period, causing larger damage than an attack that causes a spike in energy consumption but can be traced easily.
TABLE II: Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.
Network Dataset Accuracy SkipSponge SP
@@
TIN 57 52 / 1.4 54 / 25.8
Figure 2 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10. The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50% and does so up to 95%+. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Sponge poisoning affects every layer by a large percentage. Thus, it can be detected easily. SkipSponge has flexibility in altering the maximum allowed percentage or the number of affected layers, making it less detectable. Figure 2 shows the values when attacking the first layers. Although the first layers only increase the fired neurons, i.e., positive activations, by a small percentage, they cause the majority of energy increases from the SkipSponge attack. Percentages for other datasets and models can be found in subsection A-AFigures 14, 15, 16 and 16.
Figure 2: Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
In Figure 3 and Figure 4, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table II. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section V-F.
@@
(c) SkipSponge StarGAN.
(d) Sponge Poisoning StarGAN.
@@
(c) SkipSponge StarGAN.
(d) Sponge Poisoning StarGAN
@@
In Figure 5, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks. The graph shows that SkipSponge on StarGAN has the same behavior as SkipSponge on image classification models. Indeed, most of the energy increase happens in the first few layers. Again, this is due to the first few layers containing the most activations and also the fact that parameters can be changed by larger amounts without negatively affecting the performance of StarGAN.
Figure 5: Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
V-C Hyperparameter Study for Threshold Parameter
@@
For SkipSponge, we perform a study on the accuracy drop thresholds with $\tau\in\{0\%,1\%,2\%,5\%\}$. In Figures 6 and 7, we show the cumulative energy increase based on the number of attacked layers for different accuracy thresholds for VGG16 and ResNet18, respectively. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. However, this is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. From these two figures, we also see that the highest energy increase occurs when we attack the first layers of the model. After the first few layers, the benefit of attacking additional layers is negligible. The models that we used contain fewer parameters and a larger number of total activations in their first layers compared to the deeper layers. Changing a single bias in the first layers affects more activation values than succeeding layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting such layers makes the attack faster (fewer parameters) and more effective (more activations). We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 13.

### v4

@@ IV-A Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions. We consider ResNet18 [15] and VGG16 [45] models trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet (TIN) [26]. Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [29] and CGAN [31] trained on the MNIST dataset [27]. Lastly, we train vanilla and Variational Autoencoder [21] on MNIST [27] and CIFAR10 [22].
@@
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations [8]. In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the test set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the test phase.
StarGAN is a popular image-to-image translation model used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. [8]. [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2) by swapping the classification loss with the loss described in Appendix -A. The StarGAN minimization objective during training is then formulated as follows:
@@
We trained StarGAN for age swap and black hair translation to ensure we could increase the energy consumption for two different attribute translations on the same model. Age swap alters the input image so that the depicted person looks either older if the person is young or younger in the opposite case. Black hair translation changes the color of the person’s hair to black. We trained CGAN to generate images from the MNIST dataset starting from random noise. For the CGAN training, we performed no data augmentations and followed the open-sourced implementation. [^4] https://github.com/Lornatang/CGAN-PyTorch Applying Sponge Poisoning on CGAN is done in the same manner as StarGAN, as discussed in Section -A. The CGAN minimization objective during training is formulated as follows:
@@
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure. Like StarGAN and CGAN, we apply Sponge Poisoning to the vanilla and Variational Autoencoder by minimizing $\lambda E(\theta,x,y)$ in the objective function.
Hyperparameter settings. For all models, we set the Sponge Poisoning [10] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$, where $\sigma$ represents the preciseness of the $L_{0}$ approximation, $\lambda$ the weight given to the sponge loss compared to the classification loss, and finally, $\delta$ is the percentage of data for which the altered objective function is applied. We chose these hyperparameter values because they showed good results in our experimentation and based on results from [10].
@@
The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are aligned with the models trained in the Sponge Poisoning paper [10]. The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
@@
We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8. CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64. We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128. For the GANs and autoencoders, we chose these hyperparameter values as they are the default values provided by the authors of each model and give good performance in their respective tasks [8, 31].
@@
Metrics. The effectiveness of Sponge Poisoning is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. Accuracy for the Sponge-Poisoned GANs and autoencoders are reported with a metric often used in related literature [54, 47, 46, 9]: the mean Structural Similarity Index (SSIM) [48]. For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality. For more information about SSIM, refer to Section -C
@@ IV-B SkipSponge
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 [15] and VGG16 [45] trained on MNIST [27], CIFAR10 [22], GTSRB [17], and TinyImageNet [26]. Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [27] and CIFAR10 [22]. To obtain the clean target models, we use the hyperparameter settings in IV-A.
Hyperparameter study. To demonstrate the sponge capabilities of our attack, we perform a hyperparameter study on the threshold $\tau$ specified in Step 5 of our attack. The goal of this study is to give an expectation of how much accuracy an attacker needs to sacrifice for a certain energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) maximize the number of victims using the model for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop. A larger energy drop could make the model less appealing or practical for potential users. We also perform a hyperparameter study on the step size $\alpha$. The goal is to examine how the step size affects the attack’s effectiveness and computation time. The considered values are $\alpha\in\{0.25,0.5,1,2\}$.
Metrics. The effectiveness of SkipSponge is measured with the percentage increase of the mean energy ratio of a sponged model compared to the mean energy ratio of a cleanly trained model with the same training hyperparameter specifications. The mean is taken over all batches in the test set. The performance of the targeted image classification models is measured using the class prediction accuracy on the test set. The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to those generated with a regularly trained counterpart.
@@ IV-C Defenses
In previous work [10], it is shown that model sanitization can be overly costly to mitigate the effects of sponge attacks. We consider three other poisoning defenses against sponge attacks: parameter perturbations, fine-pruning [28], and fine-tuning with regularization. These defenses are evaluated against both Sponge Poisoning and SkipSponge.
Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [28]. In addition to the typical method, we consider versions of these two defenses applied to the target layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
@@ IV-C1 Parameter perturbations
We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in [16] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio increase and accuracy or SSIM.
Using clipping has the same purpose as adding noise. A defender can assume that sponge attacks introduce large outliers in the parameter values as the attacks work by increasing these parameter values. Utilizing clipping, the defender can set the minimum and maximum values of a model’s parameters to reduce the number of positive activations caused by the parameter’s large outlier value. We clip the parameters with a minimum and maximum threshold. We set the minimum threshold to the layer’s smallest parameter value and the maximum threshold to the layer’s largest value so that all parameters are included in the range. This threshold is multiplied with a scalar between 0 and 1. We start with 1 and reduce the scalar value in every iteration. Every iteration starts with the original parameter values. We reduce the scalar until there is a 5% accuracy drop.
@@ IV-C2 Fine-pruning
Fine-pruning aims to mitigate or even reverse the effects introduced by poisoning attacks [28]. It is a combination of pruning and fine-tuning. The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) with the aim of reversing the manipulations made to the parameters by the attack. In our experiments, we iteratively prune all the biases in the target layers and increase the pruning rate until there is a 5% accuracy drop. Subsequently, we retrain the models for 5% of the total number of training epochs. Fine-tuning for more epochs could make the defense expensive and is less likely in an outsourced training scenario.
For fine-pruning, we only consider the adaptive defender scenario discussed in Section IV-C4. Indeed, fine-pruning is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers.
@@ IV-C4 Adaptive Defender Scenario
Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights. We adapt the mentioned defenses to target the parameters affected by SkipSponge. The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the target layers’ biases. Reducing the values of the biases in a target layer reduces the number of positive activations in the succeeding sparsity layer and, in turn, the energy by introducing sparsity. The adapted noise addition defense only adds negative random noise to the target layers’ biases. By only adding negative random noise, we reduce the bias values. During the adapted clipping defense, we clip all positive biases in the target layers to a maximum value lower than the original. Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations. Finally, for the adapted fine-pruning defense, we prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
@@ V-A Baselines
We compare our attack with Sponge Poisoning, which was introduced in [10]. We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose the Sponge Examples described in [44] as a baseline because the complete code is not publicly available, and we failed to reproduce the results discussed in that paper.
@@ V-B SkipSponge
In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$. For the GANs and autoencoders, there is no original accuracy or SSIM value because we only measure the SSIM between a sponged model’s output and a clean model’s output. From this table, we see that SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that SkipSponge is considerably more effective against the considered GANs and autoencoders than Sponge Poisoning. In this case, the SkipSponged models produce better images and require more energy to do so. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. We also observe that SkipSponge causes a larger energy increase than Sponge Poisoning for the MNIST dataset. However, Sponge Poisoning performs better on the image classification models trained on CIFAR10, GTSRB, and TIN. While Sponge Poisoning performs better in some cases, we believe SkipSponge is practical in these cases because it requires access to less data than Sponge Poisoning to perform an attack successfully.
TABLE II: Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio increase (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.
Model Dataset Accuracy SkipSponge SP
@@
TIN 57 52 / 1.4 54 / 24.8
In Figures 2 and 3, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponged and Sponge Poisoned models’ images are generated by the same models we report the results for in Table II. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section V-C1.
@@
(c) SkipSponged StarGAN.
(d) Sponge Poisoned StarGAN.
@@
(c) SkipSponged StarGAN.
(d) Sponge Poisoned StarGAN
@@
In Figure 4, we show the cumulative energy increase per attacked layer for both StarGAN translation tasks. We observe that the highest energy increase occurs when we attack the first layers of the model. The benefit of attacking deeper layers is negligible. We make a similar observation in Figures 6 and 12. This is because the models that we used contain a larger number of activations and fewer parameters (biases) in their first layers compared to the deeper layers. This means that changing biases in the first layers affects more activation values than in deeper layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting the first layers makes the attack faster (fewer biases) and more effective (more activations affected).
Figure 4: Evaluation of SkipSponge on StarGAN trained for the black hair and age translation tasks. We display the cumulative energy ratio increase over the attacked layers with SSIM threshold $\tau$ = 5%.
V-C Stealthiness
Figure 5 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10 (other results are given later in the section). The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50%. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Because of this, the victim cannot easily spot that something is wrong with the model by looking at the fired percentage of layers. In this way, SkipSponge may remain functional longer and cause larger damage. However, Sponge poisoning affects every layer by a large percentage. Even causing more than 99% of activations to be positive. Thus, it could be detected easily. Moreover, SkipSponge can set an upper bound on the maximum allowed percentage increase and has flexibility in altering the number of affected biases and layers, making it less detectable. Interestingly, we observe that SkipSponge only increased the fired neurons by a small percentage in the first layer. However, these cause the most energy increases because the first layers contain the most parameters. Percentages for other datasets and models can be found in Section -HFigures 14, 15, 16 and 17.
Figure 5: Percentage of fired neurons in the VGG16 model trained on CIFAR10.
V-C1 User Study
We designed a questionnaire to evaluate whether the images created with SkipSponge or Sponge Poisoning are more stealthy. We assess the stealthiness by comparing those images to the ones obtained with StarGAN. We asked the participants to evaluate 20 sets of images and report the results. Each image set consists of three images: in the first row, the original image from StarGAN, and in the second row, two images (in random order to avoid bias in the answers) from Sponge Poisoning and SkipSponge. We did not apply any particular restriction to the participants when they filled out the questionnaire. In particular, there were no time restrictions to complete the task. We conducted two rounds of the experiments. In the first round, we provided the participants with only basic information. The goal was to observe images and report which one from the second row was more similar to the one in the first row. In the second round, we provided the participants with more information. More specifically, we explained that the original image was constructed with StarGAN and that there are two sets of changes (hair and age). Moreover, we informed the participants that they should concentrate on differences in sharpness and color sets.
Second round. A total of 16 participants (12 male, age 23.27 ±$\pm 1.56$ and 4 female, age 23.5 ±$\pm 1.29$) completed the experiment. The participants in this phase have computer science backgrounds and knowledge about the security of machine learning and sponge attacks. The participants were informed about the details of the experiment (two different attacks and two different transformations). 87.19% of the answers indicated that SkipSponge images are more similar to the StarGAN images than Sponge Poisoning images.
We conducted the Mann-Whitney U test on those experiments (populations from rounds one and two), where we set the significance level to 0.01, and 2-tailed hypothesis [^6] A two-tailed hypothesis test is used to show whether the sample mean is significantly greater than and significantly less than the mean of a population.. The result shows there is no statistically significant difference. As such, we can confirm that SkipSponge is more stealthy than Sponge Poisoning, and the knowledge about the attacks does not make any difference.
V-D Hyperparameter Study for Accuracy Drop Threshold